OpenStudy (anonymous):

The height of a triangle is 4 centimeters less than the base. The area of the triangle is 58.5 square centimeters. Find the length of the base and height.

OpenStudy (lgbasallote):

h = b-4 b=b a = 58.5 = b/2h substitute the values...

OpenStudy (callisto):

Let b be the base of the triangle From : The height of a triangle is 4 centimeters less than the base , we know that h = b-4 Area of triangle = (1/2)bh = (1/2)b(b-4) Since area is given to be 58.5 square centimeters Therefore 58.5 = (1/2)(b)(b-4) 117 = b^2 -4b Can you solve it from here?

OpenStudy (anonymous):

working on it. thanks